OpenStudy (valerie.m):

how do i solve 3(y-7)^2=-12

TheSmartOne (thesmartone):

Are you sure that it is equal to -12? Because you won't have any real solutions if so. the square of any number will always come out positive. Unless we're solving for imaginary solutions. Divide both sides by 3 Take the square root of both sides. Remember when you take the sqrt, you will get a plus or minus. Sqrt(16) = plus or minus 4. And then add 7 on both sides. :P

OpenStudy (austinl):

I believe this problem is dealing with complex numbers. I got y=7+-i (y=7+i, y=7-i).